OpenStudy (anonymous):

52 - 36 + 85 * 68 + 5 I know how to do it but every answer I come up with is not on the paper :( The answer I got was 5801

OpenStudy (anonymous):

\(85 \times 68\) = ??

OpenStudy (anonymous):

Tell the answer buddy.. :)

OpenStudy (anonymous):

5780

OpenStudy (mokeira):

no....this is how you do it BODMAS Start with multiplication \[52-36+(85\times68)+5=52-36+5780+5\] Next step, add \[(52+5780+5)-36\] Try that now and tell me what you get

OpenStudy (anonymous):

Now add 5 + 52 to it..

OpenStudy (anonymous):

Well, you are doing right...

OpenStudy (anonymous):

Are you given with any answer choices??

OpenStudy (anonymous):

Yes the answer choices are A. 6873 B. 3876 C. 85 D. none of the above

OpenStudy (anonymous):

But I get none of the above for every question so I know something isn't right

OpenStudy (anonymous):

What man!!! You are given with D also.. Just choose that.. :P

OpenStudy (anonymous):

for every question?

OpenStudy (anonymous):

Not for every.. Here, as your answer is not in any of the choices, but D is saying that none of the above, so you will opt that..

OpenStudy (anonymous):

Mostly you will get your answer, it is a rare case you find your answer in None Of the Above..

OpenStudy (anonymous):

Alright thanks

OpenStudy (anonymous):

Welcome.. :)

OpenStudy (anonymous):

By the way Mokeira I added all that and it was the same answer

OpenStudy (mokeira):

\[(52+5780+5)-36\] should give you \[5837-36=5801\] and it is not in the choices..so , none of the above

OpenStudy (mokeira):

@Fosterthepeople